OpenStudy (anonymous):

find the LCM 15y^2+60y and 5y^2+35y+60

jimthompson5910 (jim_thompson5910):

15y^2+60y = 15y(y+4) 5y^2+35y+60 = 5(y^2+7y+12) = 5(y+4)(y+3) Now multiply the unique factors that occur the most to get 15y(y+4)(y+3) So the LCM is 15y(y+4)(y+3)